# Many-Body Localization in the Age of Classical Computing

### v2

@@ VIII.1.3 Non-equilibrium steady states in open systems
Instead of probing the unitary dynamics, an alternative way to study coherent transport in interacting many-body systems is to mimic an experimental scenario in which reservoirs of a conserved density are attached to two ends of a 1D chain. In that case, the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation [243, 375] describes the evolution of system’s density matrix,
$\frac{d\rho}{dt}=i[\rho,\hat{H}]+\frac{1}{4}\sum_{k=1}^{4}\left([L_{k}\rho,L_{k}^{\dagger}]+[L_{k},\rho L_{k}^{\dagger}]\right)\equiv\hat{\cal L}(\rho),$ (66)
where $\hat{H}$ is the Hamiltonian of the considered many-body system, and the Lindblad operators $L_{k}$ effectively account for driving by the Markovian baths attached to the end of the chain. The density matrix $\rho(t)$, initialized in a generic initial state $\rho(0)$, converges, at a sufficiently long time, to a non-equilibrium steady state (NESS) $\rho_{\infty}=\lim_{t\to\infty}{\exp{(\hat{\cal L}t)}}\rho(0)$. Keeping track of the full density matrix $\rho(t)$ requires storing $\mathcal{N}^{2}$ complex numbers, much larger than $\mathcal{N}$ numbers required to store a pure state |$\ket{\psi}$. For that reason, the exact numerical simulation of the GKSL equation (66) is feasible on present-day computers only for small systems $L\lessapprox 14$, unless the quantum jumps method to simulate (66) is employed [150].
@@
To probe spin transport in the disordered XXZ spin chain (7), the Lindblad operators in (66) can be chosen as $L_{1}=\frac{1}{2}\sqrt{1+\mu}\,S^{+}_{1},L_{2}=\frac{1}{2}\sqrt{1-\mu}\,S^{-}_{1}$, and $L_{3}=\frac{1}{2}\sqrt{1-\mu}\,S^{+}_{L},L_{4}=\frac{1}{2}\sqrt{1+\mu}\,S^{-}_{L}$. An asymmetry in driving between the two ends, i.e., $\mu\neq 0$, induces a non-vanishing value of the spin current in the NESS $j_{l}=\mathrm{tr}\left(\hat{j}_{l}\rho_{\infty}\right)$, where the spin current operator is defined as
$\hat{j}_{l}\equiv S^{x}_{l}S^{y}_{l+1}-S^{y}_{l}S^{x}_{l+1}.$ (67)
The NESS obtained in this setting has a uniform spin current profile $j_{l}\equiv j$, and the value of the spin current scales as
@@
where the exponent $\gamma$ is linked with the dynamical exponent $z$ via $\gamma=z-1$ [388]. Deep in the ergodic regime of the disordered XXZ spin chain at $0.50<W<$ 0.5, by examining large systems comprising up to $L=400$ sites, the MPS simulations of the NESS [698] find a clear diffusive behavior $j\propto L^{-1}$. This conclusion aligns with the general expectation of diffusive behavior in non-integrable ergodic systems.
With increasing the disorder strength, but still in the ergodic regime at $20.5<W<$ 2, Refs. [698, 568] report a subdiffusive behavior $j\propto L^{-\gamma}$ with $\gamma>1$. Being limited by the decay of the slowest mode, increasing $W$ therefore makes calculating the NESS $\rho_{\infty}$ more challenging, even assuming the operator entanglement does not significantly grow in this range of $W$. The slowdown of the dynamics increases the evolution times needed to converge to the NESS solution, leading to a larger amount of the operator entanglement entropy in $\rho_{\infty}$. Consequently, at $W=2$, the MPS method employed in [698] provides access to results for relatively small system sizes up to $L\leq 20$, accessible also to the ED. Moreover, already at $W=0.65$, a small but noticeable trend for a decrease of the value of $\gamma$ towards unity, i.e., towards the diffusive transport, is visible when the system size $L$ is increased. Thus, the presently available numerical results do not allow for clear conclusions about those disorder strengths, similar to the investigations of the unitary time dynamics. Finally, we note that the spin transport properties on which we focus here may differ from the energy transport in disordered models [414].
The importance of finite size effects on the putative transition between diffusive and sub-diffusive dynamics in disordered XXZ spin chains is further exemplified by a random dephasing model studied in [644]. The random dephasing model is a non-interacting 1D Anderson insulator (5) with randomly positioned on-site dephasing terms.
@@ VIII.1.4 Conclusion
While there is clear evidence for diffusive transport deep in the ergodic regime of the disordered XXZ spin provided by several complementary methods, the transport properties at larger disorder strength $W$ remain unclear. The increase of $W$ leads to a slowdown of dynamics, hindering the assessment of the class of transport both in the NESS approach to the disordered XXZ spin chain and in the direct examinations of the unitary dynamics of the model. Investigations of the dynamical spin conductivity run into a similar barrier when the extracted diffusion constant decreases and starts to be comparable to the many-body level spacing.
In a broader perspective, the outlined problems leave the question about the relation of transport and thermalization in disordered spin chains at least partially open. This question is of particular conceptual relevance as both diffusion and subdiffusion may be expected, on the grounds of classical random walks considerations (see App. XIV), to be consistent with relaxation to thermal equilibrium. The uncertainty of the position of the putative MBL transition makes the question about the fate of the ergodic regime of disordered spin chains even more exciting.
@@ VIII.2 Transport at large disorder strengths
The difficulty of studying coherent transport in disordered spin chains can be directly linked to the slowdown of the dynamics at increasing disorder $W$. To show that [462] assume that the disordered XXZ spin chain is initialized in a weakly polarized domain wall state $\rho(0)\propto\prod_{k=1}^{L/2}\left(1+\mu S^{z}_{k}\right)\otimes\prod_{k=L/2+1}^{L}\left(1-\mu S^{z}_{k}\right)$, where $\mu\ll 1$ describes the initial polarization of the system. The domain wall decay is quantified by $\Delta S(t)=\mu\frac{L}{2}-\sum_{k=1}^{L/2}\mathrm{tr}\left(\rho(t)S^{z}_{k}\right)$, which quantifies how many particles (when the XXZ spin-1/2 chain is expressed in terms of spinless fermions) have moved after a time $t$.
@@
At small disorder strength $W=0.5$, a power-law growth $\Delta S(t)\propto t^{1/z}$ with the dynamical exponent $z=2$ is observed, consistently with the diffusion in the system. While the MPS simulation employed in [462] does not suffer from the effects of finite system size, it can access only relatively small times $t<100$. At larger disorder strengths $W$, the exponent $z$, quantifying the behavior of $\Delta S(t)$, systematically decreases with increasing time $t$. For instance, at $W=2$ the obtained exponent is $z\approx 3.8,3.3,3.1$ respectively for the fitting windows $t\in[20,40],[40,60],[60,80$]. Hence, the observed subdiffusion may be a transient effect, consistent with the conclusions of the earlier studies.
@@
An increase in the disorder strength leads to a dramatic growth of the time $t_{N}$ at which $N$ particles cross the position of the domain wall. The time $t_{1}$, after which a single particle is transported over the interface between the domains, increases exponentially with disorder strength, $t_{1}\propto e^{aW}$, for $W<3$, and the increase becomes even faster at larger disorder strengths. To make substantial statements about the nature of the transport in the system, the number of the particles transferred between the subsystems should scale with system size $L$, e.g., proportionally to $L$. The simulations of [462], based on extrapolation of the short-time behavior, suggest that the transport time $t_{N}$ exceeds the Heisenberg time $t_{H}$ with increasing disorder strength. For instance, the time required to transport $N=L/5=4$ particles exceeds the Heisenberg time for $L=20$ at $W=3$. This implies that attempts to understand the transport properties at $W=3$ must be performed at system sizes significantly exceeding $L=20$ and involve probing the dynamics at much larger time scales, e.g., at times larger than the Heisenberg time for $L=20$. The further slowdown of the dynamics at larger disorder strength $W$ results in progressively stronger conditions on the required system sizes and time scales.
@@
Several works  [58, 48, 270] investigated the behavior of dynamical spin conductivity (63) in disordered spin chains at disorder strengths exceeding the boundary of the ergodic regime, $W>W^{T}(L)$. These studies unravel that the exponential decrease of the DC conductivity $\sigma_{0}\propto e^{-aW}$ persists to disorder strengths exceeding the onset of the MBL regime, $W>W^{*}(L)$. This behavior is another manifestation of the abrupt slowdown of the dynamics with increasing disorder strength. However, the validity of the assumptions about the independence of the broadening of the Dirac delta functions in (63), required for the determination of $\sigma_{0}$ in this regime of disorder strengths, may not be taken for granted [503, 388].
@@
The results discussed in this Section show that the persistent slow dynamics at significant disorder strengths prevent not only reaching conclusions about the MBL phase but also about the character of transport in disorder spin chains when the disorder strength increases.
@@ IX.1 Rare thermal bubbles
The uncorrelated nature of the on-site magnetic fields assumed in the disordered spin chains typically studied in the context of MBL, e.g., in the XXZ model (7) or in the $J_{1}$-$J_{2}$ model (27), necessarily leads to the appearance of inhomogeneities in the system. The probability $p_{\ell}$ that the on-site potentials $h_{i}$ fulfill $|h_{i}|<W_{0}$ on $\ell$ consecutive sites, where $W_{0}$ is a constant smaller than the disorder strength $W$, is given by $p_{\ell}=(W_{0}/W)^{\ell}=e^{-\ell\log(W/W_{0})}$. The probability $p_{\ell}$ is exponentially suppressed in $\ell$, and the regions with locally weaker disorder strength are rare. Nevertheless, such rare regions of arbitrary size $\ell$ exist in the 1D chain with probability $1$ in the thermodynamic limit $L\to\infty$, and understanding their effects on the behavior of disordered spin chains is of vital importance.
@@ IX.1.1 Quantum avalanches
Figure 14: The avalanche scenario. Subsystem $A$ is subject to strong disorder and is in the MBL regime. The subsystem $B$ is weakly disordered and in the ergodic regime. The two subsystems interact via boundary term $H_{AB}$, which may induce a gradual thermalization of the subsystem $A$. Reprinted and adapted from [634].
The avalanche mechanism was introduced by [158], who proposed that the rare regions of anomalously weak disorder in the many-body system may become ergodic "bubbles" that destabilize the surrounding MBL regions. To describe this mechanism, let us consider a disordered spin chain in which a rare region with small local disorder strength $W_{B}$ and size $L_{B}$ is formed [^34] The following exposition of the avalanche mechanism is similar to descriptions of [493] and [634].. The weakly disordered subsystem $B$ is in contact with subsystem $A$ of size $L_{A}$ subject to disorder strength $W_{A}$, see Fig. 14. The Hamiltonian of the disordered spin chain, e.g., the XXZ model (7), corresponding to this scenario can be expressed as
$H=H_{A}+H_{B}+H_{AB},$ (69)
where $H_{A}$ describes the MBL subsystem $A$, $H_{B}$ is the Hamiltonian of the ergodic bubble, and $H_{AB}$ involves local operators with support at the boundary between the subsystems $A$ and $B$.
@@
    Matrix elements $\braket{m_{B}}{O}{n_{B}}$ of operators $O$ with support in the subsystem $B$, i.e., in the ergodic bubble, between the eigenstates |$\ket{m_{B}}$ of $H_{B}$ ($H_{B}\ket{m_{B}}=E_{B}\ket{m_{B}}$) follow the ETH (3), i.e.,
    $\bra{m_{B}}O\ket{n_{B}}=O(\bar{E})\delta_{n_{B}m_{B}}$
@@
    where $e^{-S(\bar{E})/2}=\rho_{B}^{-1/2}$, with $\rho_{B}$ the density of many-body states in the subsystem $B$.
@@
    The Hamiltonian of the MBL subsystem $A$ can be expressed in terms of the set of LIOM operators, $\{I_{\alpha}\}$, $\alpha=1,\dots L_{A}$. Each of the LIOMs can be expanded in the microscopic spin operators $(\hat{S}^{x}_{j},\hat{S}^{y}_{j},\hat{S}^{z}_{j})$. The norm of the terms in the expansion decreases exponentially with characteristic lengthscale $\xi$ with the distance from the center of localization of the LIOM $I_{\alpha}$. For simplicity, one may consider replacing $I_{\alpha}$ with a simple spin operator, such as $I_{\alpha}\to S_{j}^{z}$.
@@
    The coupling $H_{AB}$ involves the spin operators $(\hat{S}^{x}_{j},\hat{S}^{y}_{j},\hat{S}^{z}_{j})$ supported at the interface between the subsystems $A$ and $B$. Hence, it can be rewritten as $H_{AB}=\sum_{j=1}^{L_{A}}H_{AB}^{(j)}$ with
@@
    where $O_{i}$ acts on the degrees of freedom of the ergodic subsystem $B$, $I^{-}_{j}$ is a spin lowering operator associated with LIOM $I_{j}$ centered at site $j$, and $r=|i-j|$ is the distance between the LIOM and the ergodic bubble. For simplicity, one may model $O_{i}\to S_{i}^{x}$ and $I_{j}^{-}\to S_{j}^{x}$, giving rise to the model similar to the one in Eq. (11).
@@
|$\ket{\phi_{\mathrm{in},n_{B}}}=\ket{I_{1}\dots I_{j}\dots I_{L_{A}}}\otimes\ket{n_{B}}$ (72)
with a fixed configuration of LIOMs in the subsystem $A$, $I_{j}\in\{\uparrow,\downarrow\}$, and with the subsystem $B$ initialized in its eigenstate |$\ket{n_{B}}$ with energy $E_{n_{B}}$. Within the first order perturbation theory, the interaction term $H^{(L_{A})}_{AB}$ couples the initial state |$\ket{\phi_{\mathrm{in}}}$ to the states of the form
|$\ket{\phi_{\mathrm{fin},m_{B}}}=\ket{I_{1}\dots I_{j}\dots\overline{I}_{L_{A}}}\otimes\ket{m_{B}}$ (73)
@@
|$\ket{\phi_{\mathrm{in},n_{B}}^{(1)}}=\sum_{m_{B}\neq n_{B}}\frac{\bra{\phi_{\mathrm{fin},m_{B}}}H_{AB}^{(L_{A})}\ket{\phi_{\mathrm{in},n_{B}}}}{E_{n_{B}}-E_{m_{B}}+2\epsilon_{L_{A}}}\ket{{\phi_{\mathrm{fin},m_{B}}}},$ (74)
where $2\epsilon_{L_{A}}$ is the energy associated with flipping the LIOM $I_{L_{A}}$ and $\epsilon_{L_{A}}$ is of the order of $W_{A}$. The sum (74) may be dominated by a single term with minimal energy denominator, which is typically of the order of the level spacing $\rho_{B}^{-1}$ of the subsystem $B$. The matrix element of $H_{AB}^{(L_{A})}$ in the numerator can be evaluated with the help of (71) and the ETH ansatz for subsystem $B$ (70). Hence, the contribution to the first order correction |$\ket{\phi_{\mathrm{fin},m_{B}}}$ is given by

### v3

@@ VIII.1.3 Non-equilibrium steady states in open systems
Instead of probing the unitary dynamics, an alternative way to study coherent transport in interacting many-body systems is to mimic an experimental scenario in which reservoirs of a conserved density are attached to two ends of a chain. In that case, the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation [259, 410] describes the evolution of system’s density matrix,
$\frac{d\rho}{dt}=i[\rho,\hat{H}]+\frac{1}{2}\sum_{k}\left([L_{k}\rho,L_{k}^{\dagger}]+[L_{k},\rho L_{k}^{\dagger}]\right)\equiv\hat{\cal L}(\rho),$ (66)
where $\hat{H}$ is the Hamiltonian of the considered many-body system, and the Lindblad operators $L_{k}$ effectively account for driving by the Markovian baths attached to the end of the chain. The density matrix $\rho(t)$, initialized in a generic initial state $\rho(0)$, converges, at a sufficiently long time, to a non-equilibrium steady state (NESS) $\hat{\rho}_{\infty}=\lim_{t\to\infty}{\exp{(\hat{\cal L}t)}}\rho(0)$. Keeping track of the full density matrix $\rho(t)$ requires storing $\mathcal{N}^{2}$ complex numbers, much larger than $\mathcal{N}$ numbers required to store a pure state |$\ket{\psi}$. For that reason, the exact numerical simulation of the GKSL equation (66) is feasible on present-day computers only for small systems $L\lessapprox 14$, unless the quantum jumps method to simulate (66) is employed [156].
@@
To probe spin transport in the disordered XXZ spin chain (7), the Lindblad operators in (66) can be chosen as $\hat{L}_{1}=\frac{1}{2}\sqrt{1+\mu}\,\hat{S}^{+}_{1},\hat{L}_{2}=\frac{1}{2}\sqrt{1-\mu}\,\hat{S}^{-}_{1}$, and $\hat{L}_{3}=\frac{1}{2}\sqrt{1-\mu}\,\hat{S}^{+}_{L},\hat{L}_{4}=\frac{1}{2}\sqrt{1+\mu}\,\hat{S}^{-}_{L}$. The asymmetry in driving between the two ends, i.e., $\mu\neq 0$, induces a non-vanishing value of the spin current $j_{l}=\mathrm{tr}\left(\hat{j}_{l}\hat{\rho}_{\infty}\right)$ in the NESS, where the spin current operator is defined as
$\hat{j}_{l}\equiv\hat{S}^{x}_{l}\hat{S}^{y}_{l+1}-\hat{S}^{y}_{l}\hat{S}^{x}_{l+1}.$ (67)
The NESS obtained in this setting has, away from the edges of the system, a uniform spin current profile $j_{l}\equiv j$, and the value of the spin current scales as
@@
where the exponent $\gamma$ is linked to the dynamical exponent $z$ via $\gamma=z-1$ [424]. Deep in the ergodic regime of the disordered XXZ spin chain at $0.50<W<$ 0.5, by examining large systems comprising up to $L=400$ sites, the matrix product operator simulations of the NESS [754] find a clear diffusive behavior $j\propto L^{-1}$. This conclusion aligns with the general expectation of diffusive behavior in non-integrable ergodic systems.
With increasing the disorder strength, but still in the ergodic regime at $20.5<W<$ 2, [754, 615] reported a subdiffusive behavior $j\propto L^{-\gamma}$, with $\gamma>1$. However, the dynamics exhibit a slowing down with the growth of $W$, leading to an increase of the decay time of the slowest mode, which makes calculating the NESS $\hat{\rho}_{\infty}$ more challenging, even if one assumes that the operator entanglement does not significantly grow in this range of $W$. Consequently, at $W=2$, the matrix product operator method employed in [754] provides access to results for relatively small system sizes up to $L\leq 20$, accessible also to the ED. Moreover, already at $W=0.65$, a small but noticeable trend for a decrease of the value of $\gamma$ towards unity, i.e., towards the diffusive transport, is visible when the system size $L$ is increased. Thus, the presently available numerical results do not allow for clear conclusions about those disorder strengths, similar to the investigations of the unitary time dynamics. Finally, we note that the spin transport properties on which we focus here may differ from the energy transport in disordered models [450].
The importance of finite size effects on the putative transition between diffusive and sub-diffusive dynamics in the disordered XXZ spin chains is further exemplified by a random dephasing model studied in [696]. The random dephasing model is a non-interacting 1D Anderson insulator (5) with randomly positioned on-site dephasing terms.
@@
An alternative approach to studying the dynamical properties of a quantum many-body system with Hamiltonian $\hat{H}$ is to investigate the NESS that emerge when the system is weakly coupled to a non-thermal bath [729]. In this approach, one chooses the same set of Lindblad operators $\{L^{\nu}_{i}\}$ with $\nu=1,...,\nu_{\mathrm{max}}$ ($\nu_{\mathrm{max}}\geq 1$ is a constant) for each lattice site $i$, uniformly injecting energy to the system [384, 392]. The local temperature profile of the resulting NESS $\hat{\rho}_{\infty}$, analyzed as a function of the coupling strength $\epsilon$ to the bath, can be used to quantify the dynamics of the underlying many-body Hamiltonian $\hat{H}$. A classical hydrodynamic description of an ergodic system, realized, for instance, by a spin chain, coupled to phonons and driven by light [391], shows that the local temperature variations fulfill $\delta T\propto\epsilon^{1/(2z)}$, where $z$ is the dynamical critical exponent. In contrast, one may expect that the local variation of system properties in an MBL regime will lead to a non-uniform density profile even in the weak coupling limit $\epsilon\to 0$. These intuitions were employed in the analysis of the results of tensor-network simulations of the GKSL equation with the dissipative bulk driving of strength $\epsilon$ [729], yielding
$\frac{\delta\beta}{\bar{\beta}}(\epsilon)\sim\left\{\begin{array}[]{ll}\epsilon^{1/(2z)},&\mathrm{ergodic}\,\,\mathrm{regime}\\ \frac{\delta\beta}{\bar{\beta}}\big{|}_{\epsilon\to 0}-b\,\epsilon+\mathcal{O}(\epsilon^{2}),&\mathrm{MBL}\,\,\mathrm{regime}\;,\end{array}\right.$
where $\delta\beta/\overline{\beta}$ is the inverse temperature variation. The non-analytical dependence of $\delta\beta/\overline{\beta}$ on the driving strength for $\epsilon\to 0$ enables the calculation of the dynamical critical exponent $z$ in the ergodic regime. The dynamical critical exponent $z$ increases abruptly beyond the diffusive value $z=2$ with the disorder strength $W$. At stronger disorder, a non-zero value of $\frac{\delta\beta}{\bar{\beta}}\big{|}_{\epsilon\to 0}$, extrapolated from the results at small but finite $\epsilon$, indicates the MBL regime. The MBL phase could be, strictly speaking, differentiated from the ergodic phase only in the weak-driving limit $\epsilon\to 0$ in the thermodynamic limit. The decrease of $\epsilon$ to 00 implies a slower convergence to the NESS $\hat{\rho}_{\infty}$, which constitutes the main challenge for numerical simulations, especially at larger disorder strength $W$.
@@ VIII.1.4 Conclusion
While there is clear evidence for diffusive transport deep in the ergodic regime of the disordered XXZ spin provided by several complementary methods, the transport properties at larger disorder strength $W$ remain unclear. The increase of $W$ leads to a slowdown of the dynamics, hindering the assessment of the class of transport both in the NESS approach to the disordered XXZ spin chain, as well as in the direct examinations of the unitary dynamics of the model. Investigations of the dynamical spin conductivity run into a similar barrier when the extracted diffusion constant decreases and starts to be comparable to the many-body level spacing.
In a broader perspective, the outlined problems leave the question about the relation of transport and thermalization in disordered spin chains at least partially open. This question is of particular conceptual relevance as both diffusion and subdiffusion may be expected, on the grounds of classical random walks considerations (see App. XIV), to be consistent with relaxation to thermal equilibrium. The uncertainty of the position of the putative MBL transition makes the question about the fate of the ergodic regime of the disordered spin chains even more exciting.
@@ VIII.2 Transport at large disorder strengths
The difficulty of studying coherent transport in disordered spin chains can be directly linked to the slowdown of the dynamics at increasing disorder $W$. To show that, [500] assumed that the disordered XXZ spin chain is initialized in a weakly polarized domain wall state, $\hat{\rho}(0)\propto\prod_{k=1}^{L/2}\left(1+\mu\hat{S}^{z}_{k}\right)\otimes\prod_{k=L/2+1}^{L}\left(1-\mu\hat{S}^{z}_{k}\right)$, where $\mu\ll 1$ describes the initial polarization of the system. The domain wall decay is quantified by $\Delta S(t)=\mu\frac{L}{2}-\sum_{k=1}^{L/2}\mathrm{tr}\left(\hat{\rho}(t)\hat{S}^{z}_{k}\right)$, which quantifies how many particles (when the XXZ spin-1/2 chain is expressed in terms of spinless fermions) have moved after a time $t$.
@@
At small disorder strength $W=0.5$, a power-law growth $\Delta S(t)\propto t^{1/z}$ with the dynamical exponent $z=2$ is observed, consistent with the emergence of diffusion. While the matrix product operator simulation employed in [500] does not suffer from the effects of finite system size, it can access only relatively small times $t<100$. At larger disorder strengths $W$, the exponent $z$, quantifying the behavior of $\Delta S(t)$, systematically decreases with increasing time $t$. For instance, at $W=2$ the obtained exponent is $z\approx 3.8,3.3,3.1$, respectively, for the fitting windows $t\in[20,40],[40,60],[60,80$]. Hence, the observed subdiffusion may be a transient effect, consistent with the conclusions of the earlier studies.
@@
An increase in the disorder strength leads to a dramatic growth of the time $t_{N}$ at which $N$ particles cross the position of the domain wall. The time $t_{1}$, after which a single particle is transported over the interface between the domains, increases exponentially with the disorder strength, $t_{1}\propto e^{aW}$, for $W<3$, and the increase becomes even faster at larger disorder strengths. To make substantial statements about the nature of the transport in the system, the number of the particles transferred between the subsystems should scale with system size $L$, e.g., proportionally to $L$. The simulations of [500], based on extrapolation of the short-time behavior, suggest that the transport time $t_{N}$ exceeds the Heisenberg time $t_{H}$ with increasing the disorder strength. For instance, the time required to transport $N=L/5=4$ particles exceeds the Heisenberg time for $L=20$ at $W=3$. This implies that attempts to understand the transport properties at $W=3$ must be performed at system sizes significantly exceeding $L=20$ and involve probing the dynamics at much larger time scales, e.g., at times larger than the Heisenberg time for $L=20$. The further slowdown of the dynamics at larger disorder strength $W$ results in progressively stronger conditions on the required system sizes and time scales.
@@
Several works  [61, 51, 289] investigated the behavior of dynamical spin conductivity (63) in disordered spin chains at disorder strengths exceeding the boundary of the ergodic regime, $W>W^{T}(L)$. These studies unravel that the exponential decrease of the DC conductivity, $\sigma_{0}\propto e^{-aW}$, persists to disorder strengths exceeding the onset of the MBL regime, $W>W^{*}(L)$. This behavior is another manifestation of the abrupt slowdown of the dynamics with increasing disorder strength. However, the validity of the assumptions about the independence of the broadening of the Dirac delta functions in (63), required for the determination of $\sigma_{0}$ in this regime of disorder strengths, may not be taken for granted [544, 424].
@@
The results discussed in this Section show that the persistent slow dynamics at significant disorder strengths prevent not only reaching conclusions about the MBL phase but also about the character of transport in the disordered spin chains when the disorder strength increases. While the numerical and experimental studies do not reach a definite conclusion about transport in strongly disordered many-body systems, the mathematical analysis in the disordered quantum Ising model [573] finds that the conductance of a chain of length $L$ vanishes faster than $1/L$, consistently with the dynamical critical exponent $z$ strictly larger than $2$.
@@ IX.1 Rare thermal bubbles
The uncorrelated nature of the on-site magnetic fields assumed in the disordered spin chains typically studied in the context of MBL, e.g., in the XXZ model (7) or in the $J_{1}$-$J_{2}$ model (27), necessarily leads to the appearance of inhomogeneities in the system. The probability $p_{\ell}$ that the on-site potentials $h_{i}$ fulfill $|h_{i}|<W_{0}$ on $\ell$ consecutive sites, where $W_{0}$ is a constant smaller than the disorder strength $W$, is given by $p_{\ell}=(W_{0}/W)^{\ell}=e^{-\ell\log(W/W_{0})}$. The probability $p_{\ell}$ is exponentially suppressed in $\ell$, and the regions with locally weaker disorder strength are rare. Nevertheless, such rare regions of arbitrary size $\ell$ exist in the chain with probability $1$ in the thermodynamic limit $L\to\infty$, and understanding their effects on the behavior of the disordered spin chains is of vital importance.
@@ IX.1.1 Quantum avalanches
Figure 14: The avalanche scenario. Subsystem $A$ is subject to strong disorder and it is in the MBL regime. The subsystem $B$ is weakly disordered and it is in the ergodic regime. The two subsystems interact via the boundary term $\hat{H}_{AB}$, which may induce a gradual thermalization of the subsystem $B$. Reprinted and adapted from [685].
The avalanche mechanism was introduced by [165], who proposed that the rare regions of anomalously weak disorder in the many-body system may become ergodic "bubbles" that destabilize the surrounding MBL regions. To describe this mechanism, let us consider a disordered spin chain in which a rare region with small local disorder strength $W_{B}$ and size $L_{B}$ is formed [^44] The following exposition of the avalanche mechanism is similar to descriptions of [533] and [685].. The weakly disordered subsystem $B$ is in contact with subsystem $A$ of size $L_{A}$ subject to disorder strength $W_{A}$, see Fig. 14. The Hamiltonian of the disordered spin chain, e.g., the XXZ model (7), corresponding to this scenario, can be expressed as
$\hat{H}=\hat{H}_{A}+\hat{H}_{B}+\hat{H}_{AB},$ (69)
where $\hat{H}_{A}$ describes the MBL subsystem $A$, $\hat{H}_{B}$ is the Hamiltonian of the ergodic bubble, and $\hat{H}_{AB}$ involves local operators with support at the boundary between the subsystems $A$ and $B$.
@@
    Matrix elements $\braket{m_{B}}{\hat{O}}{n_{B}}$ of operators $\hat{O}$ with support in the subsystem $B$, i.e., in the ergodic bubble, between the eigenstates |$\ket{m_{B}}$ of $\hat{H}_{B}$ ($\hat{H}_{B}\ket{m_{B}}=E_{m_{B}}\ket{m_{B}}$) follow the ETH (3), i.e.,
    $\bra{m_{B}}\hat{O}\ket{n_{B}}=O(\bar{E})\delta_{n_{B}m_{B}}$
@@
    where $e^{-S(\bar{E})/2}=\rho_{B}^{-1/2}$, with $\rho_{B}$ the density of many-body states in the subsystem $B$, and $\bar{E}=(E_{m_{B}}+E_{n_{B}})/2$. For simplicity, the energy $\bar{E}$ is assumed to be in the middle of the spectrum of $H_{B}$.
@@
    The Hamiltonian of the MBL subsystem $A$ can be expressed in terms of the set of LIOM operators, $\{\hat{I}_{\alpha}\}$, $\alpha=1,\dots L_{A}$. Each of the LIOMs can be expanded in the microscopic spin operators $(\hat{S}^{x}_{j},\hat{S}^{y}_{j},\hat{\hat{S}}^{z}_{j})$. The norm of the terms in the expansion decreases exponentially with characteristic lengthscale $\xi$ with the distance from the center of localization of the LIOM $\hat{I}_{\alpha}$. For simplicity, one may consider replacing $\hat{I}_{\alpha}$ with a simple spin operator, such as $\hat{I}_{\alpha}\to\hat{S}_{j}^{z}$.
@@
    The coupling $\hat{H}_{AB}$ involves the spin operators $(\hat{S}^{x}_{j},\hat{S}^{y}_{j},\hat{S}^{z}_{j})$ supported at the interface between the subsystems $A$ and $B$. Hence, it can be rewritten as $\hat{H}_{AB}=\sum_{j=1}^{L_{A}}\hat{H}_{AB}^{(j)}$, with
@@
    where $\hat{O}_{i}$ acts on the degrees of freedom of the ergodic subsystem $B$, $\hat{I}^{x}_{j}$ is a spin flip operator associated with LIOM $\hat{I}_{j}$ centered at site $j$, and $r=|i-j|$ is the distance between the LIOM and the ergodic bubble. For simplicity, one may model $\hat{O}_{i}\to\hat{S}_{i}^{x}$ and $\hat{I}_{j}^{x}\to\hat{S}_{j}^{x}$, giving rise to the model similar to the one in Eq. (11).
@@
|$\ket{\phi_{\mathrm{in},n_{B}}}=\ket{I_{1}\dots I_{j}\dots I_{L_{A}}}\otimes\ket{n_{B}}\;,$ (72)
with a fixed configuration of LIOMs in the subsystem $A$, $I_{j}\in\{\uparrow,\downarrow\}$, and with the subsystem $B$ initialized in its eigenstate |$\ket{n_{B}}$ with energy $E_{n_{B}}$. Within the first order perturbation theory, the interaction term $\hat{H}^{(L_{A})}_{AB}$ couples the initial state |$\ket{\phi_{\mathrm{in}}}$ to the states of the form
|$\ket{\phi_{\mathrm{fin},m_{B}}}=\ket{I_{1}\dots I_{j}\dots\overline{I}_{L_{A}}}\otimes\ket{m_{B}}\;,$ (73)
@@
|$\ket{\phi_{\mathrm{in},n_{B}}^{(1)}}=\sum_{m_{B}\neq n_{B}}\frac{\bra{\phi_{\mathrm{fin},m_{B}}}\hat{H}_{AB}^{(L_{A})}\ket{\phi_{\mathrm{in},n_{B}}}}{E_{n_{B}}-E_{m_{B}}+2\epsilon_{L_{A}}}\ket{{\phi_{\mathrm{fin},m_{B}}}},$ (74)
where $2\epsilon_{L_{A}}$ is the energy associated with flipping the LIOM $\hat{I}_{L_{A}}$ and $\epsilon_{L_{A}}$ is of the order of $W_{A}$. The sum (74) may be dominated by a single term with minimal energy denominator, which is typically of the order of the level spacing $\rho_{B}^{-1}$ of the subsystem $B$. The matrix element of $\hat{H}_{AB}^{(L_{A})}$ in the numerator can be evaluated with the help of (71) and the ETH ansatz for subsystem $B$ (70). Hence, the contribution to the first order correction |$\ket{\phi_{\mathrm{fin},m_{B}}}$ is given by